OpenStudy (anonymous):

2/3x = yz (solve for x)

OpenStudy (saifoo.khan):

\[x=\frac{2}{3yz}\]

OpenStudy (anonymous):

can you please explain how to get the answer ? :)

hero (hero):

is the x at the top or bottom of the fraction?

OpenStudy (saifoo.khan):

Divide yz on both sides, and multiply x on both sides.

OpenStudy (anonymous):

ok so 2/3x=yz x= yz divided by 2/3 x= yz x 3/2

OpenStudy (asnaseer):

\[\frac{2}{3x}=yz\]multiply both sides by 3x to get:\[2=3xyz\]now divide both sides by 3yz to get:\[\frac{2}{3yz}=x\]

OpenStudy (anonymous):

thank you ! :)

hero (hero):

Mintchev, go back to your problem!

OpenStudy (anonymous):

^ the x is beside the 2/3

OpenStudy (anonymous):

2/3 x = yz